OpenStudy (anonymous):

lim x➝5 ((x^2-4x-5)/(x^2-8x+15) -x)

OpenStudy (anonymous):

Have you got an estimate?

OpenStudy (anonymous):

nop i dont have

OpenStudy (anonymous):

\[\lim x➝5 ((x^2-4x-5)/(x^2-8x+15) -x)\]\[=\lim x➝5(((x+1)(x-5))/((x-3)(x-5))-x)\]\[=\lim x➝5(((x+1)/(x-3))-x)=6/2-5=-2\]

OpenStudy (anonymous):

See what you think.

OpenStudy (anonymous):

yeeaah it looks right! i start to understand limits! :) so proud of myself! thank u very much! :)

OpenStudy (anonymous):

This version might be easier to read.....\[limx➝5 (\frac{x^2-4x-5}{x^2-8x+15}-x)=limx➝5 (\frac{(x+1)(x-5)}{(x-3)(x-5)}-x)\]\[limx➝5 (\frac{(x+1)}{(x-3)}-x)=6/2-5=-2\]

OpenStudy (anonymous):

thanx thats perfect! could u tell me,do i always use this way to find a limit. or i can use something different?

OpenStudy (anonymous):

Not always..... This problem was designed to work out....

OpenStudy (anonymous):

cheers! :)